A compressible fluid has density ρ = 1000 kg/m3 + h × 20 kg/m4 , with h the depth of the fluid. What is the pressure difference between a point at the surface and a point 2 m below the surface?
Answer #1


The required pressure difference is, (AP) = pgdh = * (1000+20h)gdh = 8ſ (1000+20h )dh =(109) (100)/2) +(20){}}] = 20400 Pa